Units Unit Cost
July 1 Beginning Inventory 2000 $35
July 5 Sold 1000
July 13 Purchased 6000 37
July 17 Sold 3000
July 25 Purchased 8000 39
July 27 Sold 5000

Aircard Corporation tracks the number of units purchased and sold throughout each accounting period but applies its inventory costing method at the end of each period as if it uses a periodic inventory system. The following are the transactions for the month of July.

FIFO LIFO Weight Average Cost
Costs of Goods Available For Sale
Ending Inventory
Costs of Good Sold
Answer #1

Calculate following

FIFO LIFO Weighted average cost
Cost of goods available for sale 604000 604000 604000
Ending inventory 7000*39 = 273000 (5000*37+70000) = 255000 604000/16000*7000 = 264250
Cost of goods sold 604000-273000 = 331000 604000-255000 = 349000 604000-264250 = 339750